# Geometric Neural Network based on Phase Space for BCI-EEG decoding

### v2

@@ Abstract
Objective: The integration of Deep Learning (DL) algorithms on brain signal analysis is still in its nascent stages compared to their success in fields like Computer Vision, especially in Brain-Computer Interface (BCI), where the brain activity is decoded to control external devices without requiring muscle control. Electroencephalography (EEG) is a widely adopted choice for designing BCI systems due to its non-invasive and cost-effective nature and excellent temporal resolution. Still, it comes at the expense of limited training data, poor signal-to-noise, and a large variability across and within-subject recordings. Finally, setting up a BCI system with many electrodes takes a long time, hindering the widespread adoption of reliable DL architectures in BCIs outside research laboratories. To improve adoption, we need to improve user comfort using, for instance, reliable algorithms that operate with few electrodes. Approach: Our research aims to develop a DL algorithm that delivers effective results with a limited number of electrodes. Taking advantage of the Augmented Covariance Method with SPDNet, we propose the Phase-SPDNet architecture and analyze its performance and computational impact, as well as the interpretability of the results. The evaluation is conducted on 5-fold cross-validation, using only three electrodes positioned above the Motor Cortex. The methodology was tested on nearly 100 subjects from several open-source datasets using the Mother Of All BCI Benchmark (MOABB) framework. Main results: The results of our Phase-SPDNet demonstrate that the augmented approach combined with the SPDNet significantly outperforms all the current state-of-the-art DL architecture in MI decoding. Significance: This new architecture is explainable, with a low number of trainable parameters and a reduced carbon footprint.
@@ 1 Introduction
Brain-computer interface (BCI) technology allows direct communication between a user’s brain activity and external devices. Originally designed to help people with disabilities [1], its uses are expanding to other fields, such as rehabilitation using virtual reality [2, 3]. Different signals acquired from brain activity can be used for such a technology, but Electroencephalography (EEG) is a widely adopted choice, as it is a non-invasive, portable, and inexpensive methodology with a very good temporal resolution. Motor imagery (MI) tasks are largely investigated among the brain activities considered when designing a BCI. As the subject is asked to mentally execute a movement without actually performing it, it provides an asynchronous and internal control scheme with no requirement for muscle capability.
@@
This research focuses on developing a novel DL architecture, Phase-SPDNet , that outperforms the state-of-the-art classification when using a limited number of electrodes. Building on the Augmented Covariance Method (ACM) [7] that is an extension of the spatial covariance and the Symmetric Positive Definite (SPD) Neural Network -SPDNet [8], we study the Phase-SPDNet impact on the performance when using a reduced number of electrodes. We also conducted a comprehensive analysis of the model size, carbon emissions, and, ultimately, its explainability with respect to the standard SPDNet.
Spatial covariance, however, is not the only one that can be extracted from the EEG signal. Another potential candidate is coherence, which provides an alternative perspective on EEG signal characteristics. Historically, coherence has proven to be an inherently unstable feature, which is challenging to compute accurately, resulting in less robust results with respect to the spatial covariance. However, the use of information from both covariance and coherence has been shown to increase performance as we are considering an estimator of the interactions between brain areas/electrodes [9]. We are thus interested to study coherence in combination with SPDNet and especially to use the ACM methodology adapted for the coherence matrix.
Our methodology is tested through a 5-fold cross-validation evaluation (Within-Session), utilizing only three electrodes strategically positioned above the Motor Cortex. To validate our algorithm, we test our approach on almost 100 subjects from openly available datasets using the Mother Of All BCI Benchmark (MOABB) framework [10]. This research not only contributes to the advancement of EEG MI classification but also emphasizes the importance of developing efficient, user-friendly algorithms with a minimal environmental impact in the broader context of BCI technology. Additionally, our study places a strong emphasis on ensuring that our method is both reproducible and interpretable.
The article is organized as follows: In  section 2, we examine the current state-of-the-art in Deep Learning (DL) EEG decoding, emphasizing distinctions from our approach. section 3 provides an overview of the theoretical foundations of our model and the considered datasets. The obtained results from the Within-session evaluation are presented in  section 4. Subsequently, section 5 focuses on the method’s impact and current limitations, specifically focusing on its explainability and environmental impact. Finally, section 6 summarizes the findings of our study.
@@ 2.1 Machine Learning for EEG Decoding
In the domain of EEG decoding [11, 12], translating brain activity into meaningful data has become increasingly dependent on machine learning (ML) methods [13, 14, 6, 15, 9, 16, 17]. However, here, we diverge by focusing on deep learning techniques instead of a broad ML spectrum, moving beyond mere method comparison to introduce a novel EEG decoding approach.
@@ 2.3 Geometry Transformation in EEG data
In the context of delay embedding phase geometry transformation, Chen et al. [45] demonstrated the effectiveness of phase embedding in neural networks for one-dimensional bio-signal reconstruction, but without addressing the non-euclidean nature of the signal. Our methodology, in contrast, considers the non-euclidean geometry, thereby enriching the learning process. The studies by Carrara and Papadopoulo [7] and Zhou et al. [46] align with our approach in their use of phase delays and (cross-)covariance matrices for EEG decoding. Our work, on the other hand, differentiates by fully leveraging the capabilities of the SPD matrix in Riemannian Neural Networks, allowing for more efficient, interpretable layers and achieving superior performance within challenging scenarios with a reduced number of channels.
@@ 3 Material and Methods
In this section, we describe the EEG decoding problem, the phase space reconstruction transformation, the Riemannian manifold properties, the Symmetric Positive Definite neural network components, and the datasets and baselines.
@@ 3.1 EEG Decoding
We consider EEG signals as real-valued matrices. Let us denote each bio-signal window captured during a cognitive task (called an epoch) by $\mathbf{X}_{i}\in\mathbb{R}^{C\times T}$. Here, $C$ represents the number of channels (electrodes), $T$ is the number of time points in the window, and $i$ indexes the trial $i=1\ldots N$, where $N$ is the total number of trials. Each epoch corresponds to a specific cognitive task denoted by $y_{i}\in\mathcal{Y}$.
EEG decoding aims to construct a function $f:\mathcal{X}\rightarrow\mathcal{Y}$, which effectively maps each trial $\mathbf{X}_{i}$ to its corresponding label $y_{i}$. We can construct a neural network function represented as $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, composing a sequence of $l$ functions, formulated as $f_{\theta}=f_{\theta_{1}}\circ f_{\theta_{2}}\circ\dots\circ f_{\theta_{l}}$. In this structure, $\theta$ indicates the parameters describing each layer of the network.
@@ 3.2 Phase space reconstruction for the EEG signal
Our method applies Takens’ theorem over the minimally pre-processed epoch-cropped time series, which allows for the understanding of a system’s multi-variable dynamics through a single observable in a $\psi$-dimensional space by employing a phase embedding transformation as the initial function in our neural network [55, 56, 57]. This is made by constructing a phase space using a delay vector constructed from the original signal, thereby enabling the reconstruction of the system dynamics from these observations. Following the approach proposed by Carrara and Papadopoulo [7], it is possible to define a delay function $d_{p}(\mathbf{\mathbf{X}})=\mathbf{X}(t+p\tau,t+T-(\psi-p)\tau)$ for $p\in[0,\psi]$. This function allows the concatenation of measures on a sliding window over the observable in a delay vector defined as:
@@
where $\psi$ represents the embedding dimension, which dictates the order of magnitude of the phase space, and $\tau$ is the embedding delay, determining the temporal resolution of our analysis [58, 59]. As a result, $f_{\mathrm{delay}}(\mathbf{X})$ embodies an embedding of the original phase space into a higher-dimensional space, enabling a detailed examination of the system’s dynamics. This formulation considers the EEG signal as produced by a nonlinear dynamical system. Concerning the dimension, if the original epoch have dimension $\mathbf{X}_{i}\in\mathbb{R}^{C\times T}$ the matrix $f_{\mathrm{delay}}(\mathbf{X})\in\mathbb{R}^{C\psi\times(T-\psi\tau)}$. This function is applied on each window signal $\mathbf{X}_{i}\in\mathcal{X}$. An illustration of the process, with $\tau=2$ and $\psi=3$ is present in  Figure 1
Figure 1: Representation of the augmented procedure. The original signal comprises three electrodes, represented by the plot in blue, while the red plot represents the augmented dataset using embedding parameters $\tau=2$ and $\psi=3$.
Selecting the optimal $\tau$ and $\psi$ values can be extremely time-consuming [7, 59]. To overcome this difficulty, we use an adaption of Maximizing Derivatives On Projection (MDOP) [60, 7] method for the multi-epoch context of EEG Decoding, shown in Algorithm 1. The algorithm relies on the function $mdop\_embedding$ from the package $\mathrm{DelayEmbeddings}$ to implement the MDOP procedure. The function aims to create a nonuniform embedding that generates a vector of different lags $\bm{\tau}_{i}$.
@@
The MDOP algorithm emerged as a novel geometric method for analyzing dynamical systems, as opposed to conventional statistical and information-theoretic methods, such as mutual information and continuity statistics. Central to MDOP is optimizing an embedding to ensure the reconstructed attractor is as expanded as possible while simultaneously reducing redundancy among the delay components. The algorithm employs a recursive approach, wherein each step of the embedding cycle identifies the lag $\tau$ that yields the highest beta statistics. This chosen lag is then utilized in the subsequent reconstruction phase. This iterative process continues until the algorithm obtains satisfying embedding dimensions $\psi$, selected using the method of false nearest neighbors.
@@ 3.3.1 Symmetric Positive Definite Neural Networks Components
The bi-linear mapping level aims at creating more compact and discriminating SPD matrices just like the convolutional level, with the complication, however, that SPD matrices live in a Riemannian space. We can describe the equation as
@@
where $\mathbf{Z}_{k-1}\in\mathcal{S}_{d_{k-1}}^{++}$ and W is, for this layer, the learnable parameter. Because the output $\mathbf{Z}_{k}$ must belong to the SPD space, we require that $\textbf{W}_{k}\in\mathbb{R}^{(d_{k},d_{k-1})}$ belongs to a compact Stiefel $St(d_{k},d_{k-1})$ manifold.
@@
where $\textbf{U}_{k-1}$ and $\Sigma_{k-1}$ are not learned but obtained using eigenvalue decomposition of the previous layer, $\mathbf{Z}_{k-1}=U_{k-1}\Sigma_{k-1}U_{k-1}^{T}$. This layer does not present any trainable parameter.
@@
where again, the $\textbf{U}_{k-1}$ and $\Sigma_{k-1}$ are not learned but obtained using eigenvalue decomposition. This layer not present any trainable parameter. This logarithmic mapping is applied using the Identity as a reference matrix. Once this layer is applied, we can use the classical deep learning method in the Euclidean space as a Multi-Layer Perceptron.
Given the framework, we define our neural network, denoted as $f$, as a composition of sequential transformations:
$f_{\mathrm{Phase-SPDNet}}=f_{\mathrm{delay}}\circ f_{\mathbb{C}\text{ov}}\circ f_{\mathrm{BiMap}}\circ f_{\mathrm{ReEig}}\circ f_{\mathrm{LogEig}}\circ f_{\mathrm{MLP}},$ (9)
where each $f_{l}$ represents a specific transformation or layer within the network. The learning process is formalized as the mapping $f_{\mathrm{Phase-SPDNet}}:\mathcal{X}\rightarrow\mathcal{Y}$, which operates on the training dataset. $\theta$ denotes the set of parameters within the parameter space $\Theta$. The optimization objective is to minimize an average loss $\ell$ over the training dataset, defined as:
@@
To summarize, the resulting architecture Phase-SPDNet uses enriched SPD matrices that encapsulate a broader spectrum of information than their traditional counterparts thanks to the additional layer that implements the phase space reconstruction. In addition, such an approach allows, through the use of nonlinear systems theory, the reconstruction of a phase space that contains more information with respect to the one extracted from the original signal, allowing the use of fewer electrodes. The augmented SPD matrices require an SPDNet with a larger number of parameters, which is partly counteracted by using fewer electrodes. In particular, this representation accentuates how the input SPD matrix is modified from the initial dimension $\mathcal{S}_{C}^{++}$ to $\mathcal{S}_{C\times\psi}^{++}$.
@@
Figure 2: Overview of our approach. A. Phase space reconstruction for the EEG signal. For each trial from $\mathbf{X}$, we apply the function $f_{delay}$ to reconstruct the phase space. During each evaluation fold, we adopted the MDOP algorithm to estimate suitable function parameters for embedding dimension $\psi$ and the embedding delay $\tau$. For illustrative purposes, this figure considers 3 electrodes located over the motor cortex with parameters $\psi=2$ and $\tau=10$. B. Conversion of the phase space time-series with $f_{\mathbb{C}\text{ov}}$ for the covariance representation of the Symmetric Positive Definite Space $\mathcal{S}_{++}^{n}$, represented in green. With the parameter considered the covariance is a square matrix of dimension 6. C. The covariance feature space is then used to train the Symmetric Positive Definite Neural Network with BiMap, ReEig, LogEig, and MLP layers. Once the representation returns to the Euclidean space (red), after the LogEig, we adjust a fully connected layer to predict the associate label $y$ for each trial.
In the case of Phase-SPDNet , we set the subspace dimension of the BiMap layer at half of the input dimension [8]. However, for standard SPDNet, we maintained the subspace dimension at the same value as the original input dimension (although in this case the BiMap layer acts only as a rotation) in order to proceed with subsequent analyses concerning the explainability of the model. Anyway, diminishing the subspace dimension of half in the case of SPDNet consistently led to a reduction in performance. Note that the current implementation of SPDNet has a ReEig layer that is not scale-independent because of the parameter $\varepsilon=10^{-4}$. To enhance the architecture’s scale independence, we have introduced a standardization procedure for the raw signal, bringing every channel to a zero mean and a unit standard deviation.

### v3

@@ Abstract
Objective: The integration of Deep Learning (DL) algorithms on brain signal analysis is still in its nascent stages compared to their success in fields like Computer Vision. This is particularly true for BCI, where the brain activity is decoded to control external devices without requiring muscle control. Electroencephalography (EEG) is a widely adopted choice for designing BCI systems due to its non-invasive and cost-effective nature and excellent temporal resolution. Still, it comes at the expense of limited training data, poor signal-to-noise, and a large variability across and within-subject recordings. Finally, setting up a BCI system with many electrodes takes a long time, hindering the widespread adoption of reliable DL architectures in BCIs outside research laboratories. To improve adoption, we need to improve user comfort using, for instance, reliable algorithms that operate with few electrodes. Approach: Our research aims to develop a DL algorithm that delivers effective results with a limited number of electrodes. Taking advantage of the Augmented Covariance Method and the framework of SPDNet, we propose the Phase-SPDNet architecture and analyze its performance and the interpretability of the results. The evaluation is conducted on 5-fold cross-validation, using only three electrodes positioned above the Motor Cortex. The methodology was tested on nearly 100 subjects from several open-source datasets using the Mother Of All BCI Benchmark (MOABB) framework. Main results: The results of our Phase-SPDNet demonstrate that the augmented approach combined with the SPDNet significantly outperforms all the current state-of-the-art DL architecture in MI decoding. Significance: This new architecture is explainable and with a low number of trainable parameters.
@@ 1 Introduction
Brain-computer interface (BCI) technology allows direct communication between a user’s brain activity and external devices. Originally designed to help people with disabilities [1], its uses are expanding to other fields, such as rehabilitation and virtual reality [2, 3]. Different signals acquired from brain activity can be used for such a technology, but Electroencephalography (EEG) is a widely adopted choice, as it is a non-invasive, portable, and inexpensive methodology with a very good temporal resolution. Motor imagery (MI) tasks are largely investigated among the brain activities considered when designing a BCI. As the subject is asked to mentally execute a movement without actually performing it, it provides an asynchronous and internal control scheme with no requirement for muscle capability.
@@
This research focuses on developing a novel DL architecture, Phase-SPDNet , that outperforms the state-of-the-art classification when using a limited number of electrodes. Building on the Augmented Covariance Method (ACM) [7] that is an extension of the spatial covariance and the Symmetric Positive Definite (SPD) Neural Network -SPDNet [8], we study the Phase-SPDNet impact on the performance when using a reduced number of electrodes. We also conducted a comprehensive analysis of the model size and its explainability with respect to the standard SPDNet.
Spatial covariance, however, is not the only one that can be extracted from the EEG signal. Another potential candidate is coherence, which provides an alternative perspective on EEG signal characteristics. Historically, coherence has proven to be an inherently unstable feature, which is challenging to compute accurately, resulting in less robust results with respect to the spatial covariance. However, the use of information from both covariance and coherence has been shown to increase performance as we are considering an estimator of the interactions between brain areas/electrodes [9]. We are thus interested in studying coherence in combination with SPDNet and especially in using the ACM methodology adapted for the coherence matrix.
Our methodology is tested through a 5-fold cross-validation evaluation (Within-Session), using only three electrodes strategically positioned above the Motor Cortex. To validate our algorithm, we test our approach on almost 100 subjects from openly available datasets using the Mother Of All BCI Benchmark (MOABB) framework [10]. This research not only contributes to the advancement of EEG MI classification but also emphasizes the importance of developing efficient, user-friendly algorithms in the broader context of BCI technology. Additionally, our study places a strong emphasis on ensuring that our method is both reproducible and interpretable.
The article is organized as follows: In  section 2, we examine the current state-of-the-art in Deep Learning (DL) EEG decoding, emphasizing distinctions from our approach. section 3 provides an overview of the theoretical foundations of our model and the considered datasets. The obtained results from the Within-session evaluation are presented in  section 4. Subsequently, section 5 focuses on the method’s impact and current limitations, specifically focusing on its explainability. Finally, section 6 summarizes the findings of our study.
@@ 2.1 Machine Learning for EEG Decoding
In the domain of EEG decoding [11, 12], translating brain activity into meaningful data has become increasingly dependent on machine learning (ML) methods [13, 14, 6, 15, 9, 16, 17, 18, 19, 20]. However, here, we diverge by focusing on deep learning techniques instead of a broad ML spectrum, moving beyond mere method comparison to introduce a novel EEG decoding approach.
@@ 2.3 Geometry Transformation in EEG data
In the context of delay embedding phase geometry transformation, Chen et al. [48] demonstrated the effectiveness of phase embedding in neural networks for one-dimensional bio-signal reconstruction, but without addressing the non-euclidean nature of the signal. Our methodology, in contrast, considers the non-euclidean geometry, thereby enriching the learning process. The studies by Carrara and Papadopoulo [7] and Zhou et al. [49] align with our approach in their use of phase delays and (cross-)covariance matrices for EEG decoding. Our work, on the other hand, fully leverages the capabilities of the SPD matrix in Riemannian Neural Networks, allowing for more efficient, interpretable layers and achieving superior performance within challenging scenarios with a reduced number of channels.
@@ 3 Material and Methods
In this section, we briefly recall the EEG decoding problem, the Phase Space Reconstruction - PSR transformation, the Riemannian manifold properties, the SPDNet components and the datasets considered.
@@ 3.1 EEG Decoding
We consider EEG signals as real-valued matrices. Let us denote each bio-signal window captured during a cognitive task (called an epoch) by $\mathbf{X}_{i}\in\mathcal{X}=\mathbb{R}^{C\times T}$. Here, $C$ represents the number of channels (electrodes), $T$ is the number of time points in the window, and $i$ indexes the trial $i=1\ldots N$, where $N$ is the total number of trials. Each epoch corresponds to a specific cognitive task denoted by $y_{i}\in\mathcal{Y}$.
EEG decoding aims to construct a function $f:\mathcal{X}\rightarrow\mathcal{Y}$, which effectively maps each trial $\mathbf{X}_{i}$ to its corresponding label $y_{i}$. We can construct a neural network function represented as $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, composed of sequence of $l$ functions, formulated as $f_{\theta}=f_{\theta_{l}}\circ\dots\circ f_{\theta_{2}}\circ f_{\theta_{1}}$. In this structure, $\theta_{i}$ indicates the parameters describing each layer $i$ of the network and $\theta$ is the concatenation of all the parameters $\theta_{i}$.
@@ 3.2 Phase space reconstruction for the EEG signal
Our method applies Takens’ theorem over the minimally pre-processed epoch-cropped time series, which allows for the understanding of a system’s multi-variable dynamics through an embedding in a $\psi$-dimensional space by employing a phase embedding transformation as the initial function in our neural network [58, 59, 60]. This is made by constructing a phase space using a delay vector constructed from the original signal, thereby enabling the reconstruction of the system dynamics from these observations. Following the approach proposed by Carrara and Papadopoulo [7], it is possible to define a delay function $d_{p}(\mathbf{\mathbf{X}})=\mathbf{X}[t+p\tau,t+T-(\psi-p)\tau]$ for $p\in[0,\psi]$. This function allows the concatenation of measures on a sliding window over the observable in a delay vector defined as:
@@
where $\psi$ represents the embedding dimension, which dictates the order of magnitude of the phase space, and $\tau$ is the embedding delay [61, 62]. As a result, $f_{\mathrm{delay}}(\mathbf{X})$ embodies an embedding of the original phase space into a higher-dimensional space, enabling a detailed examination of the system’s dynamics. This formulation considers the EEG signal as produced by a nonlinear dynamical system. Concerning the dimension, if the original epoch have dimension $\mathbf{X}_{i}\in\mathbb{R}^{C\times T}$ the matrix $f_{\mathrm{delay}}(\mathbf{X})\in\mathbb{R}^{C\psi\times(T-\psi\tau)}$. This function is applied on each window signal $\mathbf{X}_{i}\in\mathcal{X}$. An illustration of the process, with $\tau=5$ and $\psi=2$ is present in  Figure 1.
Figure 1: Representation of the augmented procedure. The original signal comprises three electrodes, represented by the plot in blue, while the red plot represents the augmented dataset using embedding parameters $\tau=5$ and $\psi=2$.
3.3 Selection of Phase and Delay
Selecting the optimal $\tau$ and $\psi$ values can be extremely time-consuming [7, 62]. To ensure the algorithm operates efficiently, we adopted two distinct methodologies: one grounded in non-linear dynamical systems theory and the other employing a optimized search strategy.
3.3.1 MDOP
One first approach to overcome this difficulty was to use the Maximizing Derivatives On Projection (MDOP) method [63, 7] for the multi-epoch context of EEG Decoding, shown in Algorithm 1. The algorithm relies on the function $mdop\_embedding$ from the package $\mathrm{DelayEmbeddings}$ to implement the MDOP procedure. The function aims to create a nonuniform embedding that generates a vector of different lags $\bm{\tau}_{i}$.
@@ 3.3.1 MDOP
The MDOP algorithm emerged as a novel geometric method for analyzing dynamical systems, as opposed to conventional statistical and information-theoretic methods, such as mutual information and continuity statistics. Central to MDOP is optimizing an embedding to ensure the reconstructed attractor is as expanded as possible while simultaneously reducing redundancy among the delay components. The algorithm employs a recursive approach, wherein each step of the embedding cycle identifies the lag $\tau$ that yields the highest beta statistics. This chosen lag is then used in the subsequent reconstruction phase. This iterative process continues until the algorithm obtains satisfying embedding dimensions $\psi$, selected using the method of false nearest neighbors.
3.3.2 Searching for better phase and delay parameters based on Optuna
For hyper-parameter Optimization of the phase and delay, we employ the random search implemented in Optuna version $0.17.0$ [64]. We assume a categorical distribution for the variables, $\psi\in[1,10]$ and $\tau\in[1,10]$, considering the principles of independence variables found in [7]. We restricted the search time for each iteration to 15 minutes to ensure a fair comparison between the models. The search for the Optuna used the Bayesian Tree-structured Parzen Estimator (TPE) Sampler, as described in [65]. To prevent data leakage, we incorporated the hyper-parameter search into a nested cross-validation framework [66], integrated into the current MOABB version 1.1.
@@ 3.4.1 Symmetric Positive Definite Neural Networks Components
The bi-linear mapping level aims at creating more compact and discriminating SPD matrices just like the convolutional layer in a standard DL network, with the complication, however, that SPD matrices live in a Riemannian space. We can describe the equation of the BiMap layer as
@@
where $\mathbf{Z}_{k-1}\in\mathcal{S}_{d_{k-1}}^{++}$ and W is, for this layer, the learnable parameter. The matrix $\textbf{W}_{k}$ is a full rank matrix in order to guarantee the output is an SPD matrix. Such matrices belong to a noncompact Stiefel manifold making the optimization problem impossible due to the absence of an upper bound on the distance function. To address this, we impose an additional constraint requiring $\textbf{W}_{k}$ to be orthogonal. This places the weight matrix on a compact Stiefel manifold, making the optimization problem solvable [8].
@@
where $\textbf{U}_{k-1}$ and $\Sigma_{k-1}$ are not learned but obtained using eigenvalue decomposition of the previous layer, $\mathbf{Z}_{k-1}=U_{k-1}\Sigma_{k-1}U_{k-1}^{T}$. This layer has no trainable parameters.
@@
where again, the $\textbf{U}_{k-1}$ and $\Sigma_{k-1}$ are not learned but obtained using eigenvalue decomposition.This layer has no trainable parameters. This logarithmic mapping is applied using the Identity as a reference matrix. Once this layer is applied, we can use the classical deep learning method in the Euclidean space as a Multi-Layer Perceptron.
3.5 Phase-SPDNet
Given the framework, we define our neural network, denoted as $f_{\mathrm{Phase-SPDNet}}$, as a composition of sequential transformations:
$f_{\mathrm{Phase-SPDNet}}=f_{\mathrm{MLP}}\circ f_{\mathrm{LogEig}}\circ f_{\mathrm{ReEig}}\circ f_{\mathrm{BiMap}}\circ f_{\mathbb{C}\text{ov}}\circ f_{\mathrm{delay}}(\mathbf{X}),$ (9)
where each $f_{l}$ represents a specific transformation or layer within the network. The learning process is formalized as the mapping $f_{\mathrm{Phase-SPDNet}}:\mathcal{X}\rightarrow\mathcal{Y}$, which operates on the training dataset. $\theta$ denotes the set of parameters within the parameter space $\Theta$, specifically the one of the BiMap and the MLP layer. The optimization objective is to minimize an average loss $\ell$ over the training dataset, defined as:
@@ 3.5 Phase-SPDNet
To summarize, the resulting architecture Phase-SPDNet uses enriched SPD matrices that encapsulate a broader spectrum of information than their traditional counterparts thanks to the additional layer that implements the phase space reconstruction. Such an approach allows, through the use of nonlinear systems theory, the reconstruction of a phase space that contains more information with respect to the one extracted from the original signal, allowing the use of fewer electrodes. The augmented SPD matrices require an SPDNet with a larger number of parameters, which is partly counteracted by using fewer electrodes. In particular, this representation accentuates how the input SPD matrix is modified from the initial dimension $\mathcal{S}_{C}^{++}$ to $\mathcal{S}_{C\times\psi}^{++}$.
@@
Figure 2: Overview of our approach. a. Phase space reconstruction for the EEG signal. For each trial from $\mathbf{X}$, we apply the function $f_{delay}$ to reconstruct the phase space. During each evaluation fold, we adopted the MDOP algorithm or optimized search to estimate suitable function parameters for embedding dimension $\psi$ and delay $\tau$. These hyperparameters are marked in red in the figure. For illustrative purposes, this figure considers 3 electrodes located over the motor cortex with parameters $\psi=2$ and $\tau=10$. b. Conversion of the phase space time-series with $f_{\mathbb{C}\text{ov}}$ for the covariance representation of the SPD, represented in green. With the parameter considered, the covariance is a square matrix of dimension 6. c. The covariance feature space is then used to train the Symmetric Positive Definite Neural Network with BiMap, ReEig, LogEig, and MLP layers. Once the representation returns to the Euclidean space (red), after the LogEig, we adjust a fully connected layer to predict the associate label $y$ for each trial. The light blue color represents the layer that possesses trainable parameters.
In the case of Phase-SPDNet , we set the subspace dimension of the BiMap layer at half of the input dimension [8]. For standard SPDNet, we maintained the subspace dimension at the same value as the original input dimension (although in this case, the BiMap layer acts only as a rotation) in order to proceed with subsequent analyses concerning the explainability of the model. Anyway, diminishing the subspace dimension of half in the case of standard SPDNet consistently led to a reduction in performance. Note that the current implementation of SPDNet has a ReEig layer that is not scale-independent because of the absolute parameter $\varepsilon=10^{-4}$. To enhance the architecture’s scale independence, we have introduced a standardization procedure for the raw signal, bringing every channel to a zero mean and a unit standard deviation.